6-Complete the table: PH РОН (H)* (ОНу 6.2 11.5 10-6.3 10-112
Answer #1

pH | poH TH+ Тон- 6.2 7.8 б.31E-07 | 1.58E-08 2.50 11.5 | 3.16E-03 | 3.16E-12 , 6.30 7.70 | 100-6.3 | 2.00E-08 2.80 11.20 1.5

pH of solution = 6.2 We know pH + pOH = 14 then pOH = 14 - pH = 14 - so pOH of given solution - = 7.8 6.2000 7.80 7.80000) pO

Given poi = 11.5 We know pH + POH = 14 then pH = 14 - pOH = 14 - so pH of given solution = = 2.5000 11.5000 2.50 The relation

Given [H+] = 10^-6.3 = we know pH --log [H+] =-log [ 5.01E-07 So pH of the given solution = 6.30 ] 5.01187E-07 = 6.30 We also

Given [OH-] = 10^-11.2 = we know pOH =-log [OH-]=-log[ 6.3090E-12 ] So pOH of the given solution = 11.20 6.31E-12 = 11.2000 W